a) The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.

Answer by stanbon(48575)

a) The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively
Quantity Eq: x + y = 38
--------------------------
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a
Value Eq: 175000x + 200,000y = 7,200,000
--------------------------------------------
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
Equation:
Multiply thru the Quantity Equation to get the following set:
175000x + 175000y = 175000*38
175000x + 200,000y = 7,200,000
------------------------------------
Subtract 1st from 2nd to get:
25000y = 550,000
y = 22 (# of $200,000 houses)
----------
Substitute into x + y = 38 to solve for"x":
x + 22 = 38
x = 16 (# of $175,000 houses)
========================
Cheers,
Stan H.